Adrian457:

Complete the square to rewrite the following equation. Identify the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

Adrian457:

@Oddmoyu123 pls help

Oddmoyu123:

can i get a pic of your square

Oddmoyu123:

if so i can help

MrMudd183:

I would combine like terms first, then I would separate the x and y from the same side

Adrian457:

Sorry there is no image, that's all the info that the question gives

Adrian457:

@MrMudd183 what are the like terms that I can combine?

MrMudd183:

X and Y

Adrian457:

x^2 + y^2 and 4x-6y?

Adrian457:

??

Adrian457:

@MrMudd183 ??

dude:

Separate the like terms, yes Keep the 4 on the other side \(\color{green}{(x^2+4x+\_\_)}+\color{red}{(y^2-6y+\_\_)}=4\) You can complete the square from then First complete the square for the x Then complete the square for the y Treat them like separate equations with different variables

Adrian457:

@dude I don't exactly understand how to solve that, can you explain it to me please?

dude:

Knowing that quadratic equations are written as \(ax^2+bx+c\) Use \(\dfrac{b}{2}^2\) \((x^2+4x+\_\_)\) =\((\dfrac{4}{2})^2=> 4\) Therefore, \((x^2+4x+4)+\color{red}{(y^2-6y+\_\_)}=4+4\) [Remember to add it on both sides] Do the same for the y value

Oddmoyu123:

i should read these things all the way before i type